OpenStudy (anonymous):

Two charges attract each other with a force of 1.5 N. What will be the force if the distance between them is reduced to one-ninth of its original value?

OpenStudy (chmvijay):

do you know coulombs law

OpenStudy (anonymous):

Yes, (k)(q)(q2)/d^2

OpenStudy (chmvijay):

what is in question says is distance between them reduced to 1/9 put d=1/9 *d right and solve

OpenStudy (anonymous):

So what would I set the two equations to?

OpenStudy (chmvijay):

F is proportional to the inverse square of distance.

OpenStudy (chmvijay):

do u know the answer LOL:)

OpenStudy (anonymous):

I get answers but they seem way off...

OpenStudy (anonymous):

When I solve them, I have so many variables, so I can set a system of equations

OpenStudy (anonymous):

*So i can't set

OpenStudy (chmvijay):

OK F = 1/d^2 = consider as proportional

OpenStudy (chmvijay):

now d becomes one by ninth mean =1/ 9 *d right

OpenStudy (anonymous):

right

OpenStudy (chmvijay):

F=1/(1/9)^2 d2

OpenStudy (chmvijay):

right

OpenStudy (anonymous):

right

OpenStudy (chmvijay):

F =1/(1/81)d2

OpenStudy (chmvijay):

F=81/d^2 right

OpenStudy (anonymous):

yeah, i have that

OpenStudy (chmvijay):

F/81=1/d2

OpenStudy (chmvijay):

1.5/81 = 0.0185 N right

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

so i get: 0.0185d^2

OpenStudy (chmvijay):

hey wait if distance ids reduced means force should be larger then initial right but we are getting less force here?????

OpenStudy (chmvijay):

its like this F2=81/d^2 right

OpenStudy (anonymous):

I have that

OpenStudy (chmvijay):

F2=F1*81 =1.5*81 =121.5 N is right

OpenStudy (anonymous):

ohhh. I didnt see that you could set the \[(k \times Q1 \times Q2)\] as a constant

OpenStudy (chmvijay):